Good Afternoon, i just wanted to know how do you get 2/9 of -36 ?
Answer by jim_thompson5910(35256)

2/9 of -36 = (2/9)*(-36) = (2*(-36))/9 = -72/9 = -8

So 2/9 of -36 is -8